An empty steel container is filled with 0.450 atm of A and 0.450 atm of B. The system is allowed to reach equilibrium according to the reaction below. If Kp = 340 for this reaction, what is the equilibrium partial pressure of C?


A (g) + B (g) ⇌ C (g)


1
Expert's answer
2021-03-22T06:31:40-0400

Given data,

The partial pressure of A is .450 atm

The partial pressure of B is . 450 atm

Kp=340

Given reaction-

A (g) + B (g) ⇌ C (g)

We know,

Kp=p(C)/[p(A)*p(B)]

340=X/[(. 450-X)(. 450-X)]

340=X/(. 450-X)2

X=340×(. 450-X)2

X=340×(. 4502-2×. 450×X +X2)

By solving this equation we get-

X=306.75

Hence the equilibrium partial pressure of C is 306.75 atm